# Code: https://github.com/icey-zhang/EfficientMFD

### v1

@@ Abstract
[…]
Keywords:
Multimodal images Image Fusion Object Detection
@@ 1 Introduction
Precise and dependable object parsing plays a fundamental role in autonomous driving, and remote sensing monitoring [6]. However, merely using visible sensors may fail to accurately recognize objects in complex environments, e.g., inclement weather. Visible-infrared image fusion (VIF) [39, 55, 52, 54] as a typical common multimodal fusion task addresses the above limitation with the complementary information. Numerous multimodal image fusion approaches [15, 16, 19, 11, 8] have been quickly developed to learn the complete representation and complementary information between different modalities. In CDDFuse [51] and DIDFuse [53], a well-designed VIF network is trained first. Then object detection (OD) network is trained using the VIF network results to validate the fusion performance in the separate optimization. Thanks to the powerful capability of deep neural networks on representation learning across multimodalities, deep model-based multimodal fusion has achieved encouraging performance. However, most of them prioritize generating visually pleasing images over addressing downstream high-level visual tasks, posing an obstacle to better object parsing.
Figure 1: Comparison of EfficientMFD with existing VIF-OD task paradigms. Our method redefines the framework composited by image fusion and object detection sub-network in a parallel structure and it can be optimized in one stage joint way which is synchronous. We improve it with a phylogeny tree design for object-region-pixel feature modeling and task alignment to eliminate the task optimization barrier. Dashed arrows indicate cutting off gradient flow.
Recently, some studies have attempted to design a joint learning method by combining the fusion network and high-level tasks e.g. object detection [38] and segmentation [20, 40]. Multimodal fusion detection (MFD) methods have become a hot research issue that VIF and OD enjoy a symbiotic relationship that benefits both parties. VIF generates more informative fused images that significantly improve the performance of OD, while OD provides VIF with rich object semantic information, as its primary objective is to locate and identify objects within the scene. MFD networks are usually designed in a cascaded structure. Joint optimization [18] adopts an OD network as a constraint to a trained VIF network, and thus forces the VIF network to generate fusion images with easily detected objects. Particularly, Zhao et. al [50] proposed a multimodal fusion detection joint learning method, integrating meta-feature embedding from object detection to further promote fusion performance by generating object semantic features through simulating meta-learning. Despite the great success achieved by previous methods as shown in Figure 1, there still remain critical issues. 1) The dominant optimization methods tend to employ a multi-step paradigm in a progressive joint fashion, which significantly undermines efficiency. 2) This optimization overemphasizes the potential of utilizing information from OD to enhance image fusion and makes it difficult to balance the parameters involved, thus being affected by the local optimal solution of individual task. Therefore, seeking unified features for each task simultaneously is still a tough issue. This paper proposed an end-to-end multimodal fusion detection algorithm, named EfficientMFD. EfficientMFD is constructed by fine-grained detail fusion network and coarse-to-fine diffusion-based object detection network, in which the intrinsic features of either one interact via synchronous joint optimization in a one-stage end-to-end manner. To enhance the compatibility of both the fine-grained detail information and semantic content, we introduce the concept of an object-region-pixel phylogenetic tree and coarse-to-fine diffusion process, aiming to mimic the visual perception needed for VIF and OD task requirements. Moreover, we establish a gradient matrix task-alignment (GMTA) to optimize shared complement components to eliminate the undesired effects of inherent optimization. Hence, the gradient can effectively converge to an optimal point with fusion detection weights. Our main contributions can be distilled as follows: (1) We believe that this is the first proposal of the efficient synchronous joint learning named EfficientMFD, which redefined both image fusion and object detection in a one-stage end-to-end manner. We show that this can substantially improve the performance of both tasks. (2) We propose a gradient matrix task-alignment method to measure and quantify the contributions of image fusion and object detection tasks, which help to better balance the stability of the training process and converge to an optimal point with fusion detection weights. (3) Extensive experiments on image fusion and object detection validate the effectiveness of the proposed approach.
@@ 2.1 Multimodal Fusion Object Detection
Due to the powerful nonlinear fitting capabilities of deep neural networks, deep learning has made significant progress in low-level vision tasks, particularly in image fusion tasks [14, 22, 15, 27, 28, 29, 31]. Early efforts [19, 23, 53, 43, 24] tended to achieve excellent fusion results by adjusting network structures or loss functions, overlooking the fact that image fusion should aim to improve the performance of downstream application tasks. Fusion images with good quality metrics may be suitable for human visual perception but may not be conducive to practical application tasks [40]. Some research has acknowledged this issue, Yuan et al[49] utilized various aligned modalities for improved oriented object detection [46, 44, 45, 47] to address the challenge of cross-modal weakly misalignment in aerial visible-infrared images. Liu et al. [18] proposing a joint learning method, pioneered the exploration of the VIF and OD combination methods. Then the optimization of the loss function for segmentation [40] and detection [38] has been validated to be effective in guiding the generation of fused images. They consider the downstream task network as an additional constraint to assist the VIF network in generating fusion results with clearer objects. Zhao et al. [50] leveraged semantic information from OD features to aid VIF and perform meta-feature embedding to generate meta-features from OD features, which are then used to guide the VIF network in learning pixel-level semantic information. Liu et al. [20] proposed multi-interactive feature learning architecture for image fusion and segmentation by enhancing fine-grained mapping of all the vital information between two tasks, so that the modality or semantic features can be fully mutual-interactive. However, OD considers the semantic understanding of objects, while VIF and segmentation primarily focus on the pixel-level relationship between image pairs. The optimization coupling of detection and fusion tasks becomes more challenging to investigate these complementary differences, from which both fusion and detection can benefit. It is worth mentioning that these visible-infrared multimodal fusion detection methods are usually designed in a cascaded structure with tedious training steps. Researchers lack the adoption of end-to-end architectures, which would enable one-step network inference to generate credible fused images and detection results through a set of network parameters.
@@ 2.2 Multi-task Learning
Multi-task learning (MTL) [3, 35] involves the simultaneous learning of multiple tasks through parameter sharing. Prior approaches involve manually crafting the architecture, wherein the bottom layers of a model are shared across tasks [2, 13]. Some approaches tailor the architecture based on task affinity [42], while others utilize techniques such as Neural Architecture Search [1, 2, 37] or routing networks [34] to autonomously discern sharing patterns and determine the architecture. An alternative method typically combines task-specific objectives into a weighted sum [12, 5, 25]. In addition, most approaches (e.g. [7, 17, 21, 30, 48]) aim to mitigate the effects of conflicting or dominating gradients. The approach of explicit gradient modulation [21, 30, 48, 36], has demonstrated superior performance in resolving conflicts between task gradients by substituting conflicting gradients with modified, non-conflicting gradients. Inspired by the above multimodal learning methods, we introduce a gradient matrix task-alignment method to align the orthogonal components contained in image fusion and object detection tasks thereby effectively eliminating the inherent optimization barrier that exists between two tasks.
@@ 3.1 Problem Formulation
VIF-OD task concentrates on generating an image that benefits emphasizing objects with superior visual perception capability. The goal of OD is to find the location and identify the class of each object in an image which can naturally provide rich semantic information along with object location information. Therefore, the motivation of OD-aware VIF is to construct a novel infrared and visible image fusion framework that can benefit from the semantic information and object location information contained in object detection. For this purpose, we suppose a pair of visible image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C_{x}}$ and infrared image $\boldsymbol{y}\in\mathbb{R}^{H\times W\times C_{y}}$. The optimization model is formulated as:
$\min_{\boldsymbol{\theta}_{\mathrm{t}}}\mathcal{L}\left(\boldsymbol{t},\mathcal{N}\left(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\theta}_{\mathrm{t}}\right)\right)$ (1)
@@
where $\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}_{\mathrm{u}}^{s}=\boldsymbol{\theta}_{\mathrm{d}}^{s}$, defined as the shared parameters for VIF and OD networks. $\boldsymbol{u}$ and $\boldsymbol{d}$ denote the fused image and detection result, which are produced by the VIF network $\Phi(\cdot)$ and OD network $\Psi(\cdot)$ with the learnable parameters $\boldsymbol{\theta}_{\mathrm{u}}$ and $\boldsymbol{\theta}_{\mathrm{d}}$. $w$ is a predefined weighting factor to balance the task training. $\mathcal{S}(\cdot)$ is a constrained term to jointly optimize the two tasks. In this paper, we regard the $\mathcal{S}(\cdot)$ as a feature learning constrained manner and achieve this goal by designing a gradient matrix task-alignment training scheme.
Figure 2: An overview of the proposed EfficientMFD framework, which consists of a backbone, nodes, and branches. The backbone is utilized to extract multimodal image features. A fine-grained fusion network (ORPPT) and diffusion-based object detection network (CFDP) are optimized by synchronous joint optimization in an end-to-end manner. Branches include one pixel feature mining module(PFMM) and $L$ region feature refined module (RFRM) to mine different-grained region representations. The gradient domain and conflict between the two tasks are eliminated by the gradient matrix task alignment.
@@ 3.2 Architecture
Our proposed EfficientMFD is designed with parallel principle, composited by image fusion and object detection sub-network. Details of the whole architecture are shown in Figure 2. A fusion network (ORPPT) and object detection network (CFDP) can sufficiently realize the granularity-aware detail information and semantic information extraction. Object-Region-Pixel Phylogenetic Tree. The important fact of that humans pay more attention to different regions from coarse to fine for object detection in object scale and image fusion in pixel scale. Inspired by a phylogenetic tree, to simulate humans to study the interactions of hierarchies under different granularity views, we construct an Object-Region-Pixel Phylogenetic Tree (ORPPT) as $\Phi(\cdot)$ to extract different features in multiple region scales. Given an image pair $\boldsymbol{x}\in\mathbb{R}^{W\times H\times C_{x}}$, $\boldsymbol{y}\in\mathbb{R}^{W\times H\times C_{y}}$, we firstly extract image features $f(\boldsymbol{x})$ and $f(\boldsymbol{y})$ by shared parallel backbone to save memory and computing resources. Then these features are added in channel dimension to obtain the final $L$ multimodal image features $\boldsymbol{o_{1}},...,\boldsymbol{o_{l}},...,\boldsymbol{o_{L}}\in\mathbb{R}^{W_{1}\times H_{1}\times C_{1}}$. The parameters in $f(\cdot)$ are shared with the OD network. Although $f(\boldsymbol{x})$ and $f(\boldsymbol{y})$ can describe the characteristics of visible and infrared modalities, they lack insights into the multi-granularity perspective. Therefore, we utilize the branches including the one pixel feature mining module (PFMM) and $L$ region feature refine module (RFRM) to mine the multiple granularities from coarse to fine. The PFMM $B_{0}$ is set the same with feature fusion block in the MetaFusion [50] with the input pair $\boldsymbol{x},\boldsymbol{y}$. We set $1,2,…,l,…L$ to denote each region branch. For branch $l$, a CNN $\varphi_{l}(\cdot)$ is firstly utilized to extract the granularity-wise feature $\varphi_{l}(\boldsymbol{o_{l}})\in\mathbb{R}^{W_{2}\times H_{2}\times C_{2}}$ at the region level. A set of learnable region prompts $\boldsymbol{R}_{l}=\left\{\boldsymbol{r}_{l,m}\in\mathbb{R}^{C_{2}}\right\}_{m=1}^{M_{l}}$ is introduced to define the $M_{l}$ different regions of granularity-wise feature, where $r_{l,m}$ denotes the $m^{\mathrm{th}}$ region prompt at branch $l$. Then the feature vector is mapped into the region mask $\boldsymbol{A}_{l}=\left\{\boldsymbol{a}_{l,m}\in\mathbb{R}^{W_{2}\times H_{2}}\right\}_{m=1}^{M_{l}}$ by conducting the dot product between the feature vector and region prompt followed by batch normalization and a ReLU activation:
@@
where $b_{l,m}(\boldsymbol{o_{l}})$ denotes the $m^{\mathrm{th}}$ region representation and $(i,j)$ denotes the spatial location. These region-level representations are further concatenated to form the observation $B_{l}(\boldsymbol{o_{l}})=\left[b_{l,1}(\boldsymbol{o_{l}}),b_{l,2}(\boldsymbol{o_{l}}),\ldots,b_{l,M_{l}}(\boldsymbol{o_{l}})\right]$ of branch $l$. These multi-grained attentions concentrate operation on the spatial location information and the extent of the regions which are similar to the task requirements. The $B_{1}$, $B_{2}$, …, $B_{L}$ are up-sampled to keep the consistent spatial size with the pixel-level fusion features $B_{0}$. Then, the region-level fusion features $B_{1},B_{2},...,B_{L}$ are assembled by addition operation followed by a Convolution with $1\times 1$ kernel + ReLU to reduce the channel numbers. Highlighted region features are extracted by a Convolution with $1\times 1$ kernel + Sigmoid and then injected into pixel-level features by multiplication and addition. Finally, five Convolutions with $3\times 3$ kernel + ReLU layers are constructed to reconstruct the fusion result $u$. Coarse-to-Fine Diffusion Process. Diffusion models, inspired by nonequilibrium thermodynamics, are a class of likelihood-based models. DiffusionDet [4] is the first neural network model to utilize the diffusion model for object detection, introducing a novel paradigm that achieves promising results compared to traditional object detection models. Combining diffusion simulation with the diffusion and recovery process of the object box, the Coarse-to-Fine Diffusion Process (CFDP) introduces it as an efficient detection head to assist fusion networks to focus more on object areas. CFDP model defines a Markovian chain of diffusion forward process by gradually adding noise to a set of bounding boxes. The forward noise process is defined as:

### v2

# E2E-MFD: Towards End-to-End Synchronous Multimodal Fusion Detection

@@ Abstract
Multimodal image fusion and object detection are crucial for autonomous driving. While current methods have advanced the fusion of texture details and semantic information, their complex training processes hinder broader applications. Addressing this challenge, we introduce E2E-MFD, a novel end-to-end algorithm for multimodal fusion detection. E2E-MFD streamlines the process, achieving high performance with a single training phase. It employs synchronous joint optimization across components to avoid suboptimal solutions tied to individual tasks. Furthermore, it implements a comprehensive optimization strategy in the gradient matrix for shared parameters, ensuring convergence to an optimal fusion detection configuration. Our extensive testing on multiple public datasets reveals E2E-MFD’s superior capabilities, showcasing not only visually appealing image fusion but also impressive detection outcomes, such as a 3.9% and 2.0% $\text{mAP}_{50}$ increase on horizontal object detection dataset M3FD and oriented object detection dataset DroneVehicle, respectively, compared to state-of-the-art approaches.
@@ 1 Introduction
Precise and reliable object parsing is critical in fields such as autonomous driving and remote sensing monitoring [1]. Relying solely on visible sensors can lead to inaccuracies in object recognition in challenging environments, like inclement weather conditions. Visible-infrared image fusion [2; 3; 4; 5] as a typical common multimodal fusion (MF) task addresses these challenges by leveraging complementary information from different modalities, leading to the rapid development of various multimodal image fusion techniques [6; 7; 8; 9; 10]. Techniques like CDDFuse [11] and DIDFuse [12] employ a two-step process where a MF network is trained initially, followed by training an object detection (OD) network with the results from the MF network to assess fusion effectiveness separately. Although deep neural networks have significantly enhanced the ability to learn representations across modalities, resulting in promising multimodal fusion outcomes, the focus has predominantly been on producing visually appealing images. This emphasis often overlooks the improvement of downstream high-level visual tasks, such as enhanced object parsing, which remains a substantial hurdle.
Figure 1: Comparison of (d) E2E-MFD with existing MF-OD task paradigms (a) Two-Stage (Separate Cascaded), (b) Two-stage (Joint Cascaded) and (c) Multi-stage (Joint Cascaded).
Recent studies have ventured into designing joint learning methods that integrate fusion networks with high-level tasks such as object detection [13] and segmentation [14; 15]. The synergy between MF and OD in Multimodal Fusion Detection (MFD) methods has emerged as a vibrant area of research. This partnership allows MF to produce richer, more informative images, enhancing OD performance, while OD contributes valuable object semantic insights to MF, aiming to accurately locate and identify objects in a scene. Typically, MFD networks adopt a cascaded design where joint optimization techniques [16] use the OD network to guide the MF network toward creating images that facilitate easier object detection. Notably, Zhao et. al [17] introduced a joint learning method for multimodal fusion detection, incorporating meta-feature embedding from OD to improve fusion by generating semantic object features through meta-learning simulation. Despite these advancements, as highlighted in Figure 1, significant challenges persist: 1) Current optimization approaches rely on a multi-step, progressively joint method, compromising efficiency; 2) These methods overly focus on leveraging OD information for fusion enhancement, leading to difficulty in parameter balancing and susceptibility to local optima of individual tasks. Therefore, the quest for a unified feature set that simultaneously caters to each task remains formidable.
In this paper, we introduce E2E-MFD, an end-to-end algorithm for multimodal fusion detection, designed to seamlessly blend detailed image fusion and object detection from coarse to fine levels. E2E-MFD facilitates the interaction of intrinsic features from both domains through synchronous joint optimization, allowing for a streamlined, one-stage process. To reconcile fine-grained details with semantic information, we propose the novel concept of an Object-Region-Pixel Phylogenetic Tree (ORPPT) coupled with a coarse-to-fine diffusion processing (CFDP) mechanism. This approach is inspired by the natural process of visual perception, tailored to meet the specific needs of MF and OD. Furthermore, we introduce a Gradient Matrix Task-Alignment (GMTA) technique to fine-tune the optimization of shared components, thereby minimizing the adverse impacts traditionally associated with inherent optimization challenges. This ensures an efficient convergence towards an optimal set of fusion detection weights, enhancing both the accuracy and efficacy of multimodal fusion detection tasks.
Our contributions in this paper are highlighted as follows: (1) We present E2E-MFD, a pioneering approach to efficient synchronous joint learning, innovatively integrating image fusion and object detection into a single-stage, end-to-end framework. This methodology significantly enhances the outcomes of both tasks. (2) We introduce a novel Gradient Matrix Task-Alignment (GMTA) technique, designed to evaluate and quantify the impacts of the image fusion and object detection tasks. This aids in optimizing the training process’s stability and ensures convergence to an optimal configuration of fusion detection weights. (3) Through comprehensive experimentation on image fusion and object detection, we demonstrate the efficacy and robustness of our proposed method.
@@ 2.1 Multimodal Fusion Object Detection
Due to the powerful nonlinear fitting capabilities of deep neural networks, deep learning has made significant progress in low-level vision tasks, particularly in image fusion tasks [18; 19; 6; 20; 21; 22; 23]. Early efforts [8; 24; 12; 25; 26] tended to achieve excellent fusion results by adjusting network structures or loss functions, overlooking the fact that image fusion should aim to improve the performance of downstream application tasks. Fusion images with good quality metrics may be suitable for human visual perception but may not be conducive to practical application tasks [15]. Some research has acknowledged this issue, Yuan et al[27] utilized various aligned modalities for improved oriented object detection [28; 29; 30; 31] to address the challenge of cross-modal weakly misalignment in aerial visible-infrared images. Liu et al. [16] proposing a joint learning method, pioneered the exploration of the MF and OD combination methods. Then the optimization of the loss function for segmentation [15] and detection [13] has been validated to be effective in guiding the generation of fused images. They consider the downstream task network as an additional constraint to assist the MF network in generating fusion results with clearer objects. Zhao et al. [17] leveraged semantic information from OD features to aid MF and perform meta-feature embedding to generate meta-features from OD features, which are then used to guide the MF network in learning pixel-level semantic information. Liu et al. [14] proposed multi-interactive feature learning architecture for image fusion and segmentation by enhancing fine-grained mapping of all the vital information between two tasks, so that the modality or semantic features can be fully mutual-interactive. However, OD considers the semantic understanding of objects, while MF and segmentation primarily focus on the pixel-level relationship between image pairs. The optimization coupling of detection and fusion tasks becomes more challenging to investigate these complementary differences, from which both fusion and detection can benefit. It is worth mentioning that these visible-infrared multimodal fusion detection methods are usually designed in a cascaded structure with tedious training steps. Researchers lack the adoption of end-to-end architectures, which would enable one-step network inference to generate credible fused images and detection results through a set of network parameters.
@@ 2.2 Multi-task Learning
Multi-task learning (MTL) [32; 33] involves the simultaneous learning of multiple tasks through parameter sharing. Prior approaches involve manually crafting the architecture, wherein the bottom layers of a model are shared across tasks [34; 35]. Some approaches tailor the architecture based on task affinity [36], while others utilize techniques such as Neural Architecture Search [37; 34; 38] or routing networks [39] to autonomously discern sharing patterns and determine the architecture. An alternative method typically combines task-specific objectives into a weighted sum [40; 41; 42]. In addition, most approaches (e.g. [43; 44; 45; 46; 47]) aim to mitigate the effects of conflicting or dominating gradients. The approach of explicit gradient modulation [45; 46; 47; 48], has demonstrated superior performance in resolving conflicts between task gradients by substituting conflicting gradients with modified, non-conflicting gradients. Inspired by the above multimodal learning methods, we introduce a Gradient Matrix Task-Alignment method to align the orthogonal components contained in image fusion and object detection tasks thereby effectively eliminating the inherent optimization barrier that exists between two tasks.
@@ 3.1 Problem Formulation
MF-OD task concentrates on generating an image that benefits emphasizing objects with superior visual perception capability. The goal of OD is to find the location and identify the class of each object in an image which can naturally provide rich semantic information along with object location information. Therefore, the motivation of OD-aware MF is to construct a novel infrared and visible image fusion framework that can benefit from the semantic information and object location information contained in object detection. For this purpose, we suppose a pair of visible image $\boldsymbol{x}\in\mathbb{R}^{H\times W\times C_{x}}$ and infrared image $\boldsymbol{y}\in\mathbb{R}^{H\times W\times C_{y}}$. The optimization model is formulated as:
$\min_{\boldsymbol{\theta}_{\mathrm{t}}}\mathcal{L}\left(\boldsymbol{t},\mathcal{N}\left(\boldsymbol{x},\boldsymbol{y};\boldsymbol{\theta}_{\mathrm{t}}\right)\right),$ (1)
@@
where $\boldsymbol{\theta}^{\star}=\boldsymbol{\theta}_{\mathrm{u}}^{s}=\boldsymbol{\theta}_{\mathrm{d}}^{s}$, defined as the shared parameters for MF and OD networks. $\boldsymbol{u}$ and $\boldsymbol{d}$ denote the fused image and detection result, which are produced by the MF network $\Phi(\cdot)$ and OD network $\Psi(\cdot)$ with the learnable parameters $\boldsymbol{\theta}_{\mathrm{u}}$ and $\boldsymbol{\theta}_{\mathrm{d}}$. $w$ is a predefined weighting factor to balance the task training. $\mathcal{S}(\cdot)$ is a constrained term to jointly optimize the two tasks. In this paper, we regard the $\mathcal{S}(\cdot)$ as a feature learning constrained manner and achieve this goal by designing a Gradient Matrix Task-Alignment training scheme.
Figure 2: An overview of the proposed E2E-MFD framework, which consists of a backbone, nodes, and branches. The backbone is utilized to extract multimodal image features. A fine-grained fusion network (ORPPT) and diffusion-based object detection network (CFDP) are optimized by synchronous joint optimization (GMTA) in an end-to-end manner.
@@ 3.2 Architecture
Our proposed E2E-MFD is designed with parallel principle, composited by image fusion and object detection sub-network. Details of the whole architecture are shown in Figure 2. A fusion network (ORPPT) and object detection network (CFDP) can sufficiently realize the granularity-aware detail information and semantic information extraction.
Object-Region-Pixel Phylogenetic Tree. The important fact of that humans pay more attention to different regions from coarse to fine for object detection in object scale and image fusion in pixel scale. Inspired by a phylogenetic tree, to simulate humans to study the interactions of hierarchies under different granularity views, we construct an Object-Region-Pixel Phylogenetic Tree (ORPPT) as $\Phi(\cdot)$ to extract different features in multiple region scales. Given an image pair $\boldsymbol{x}\in\mathbb{R}^{W\times H\times C_{x}}$, $\boldsymbol{y}\in\mathbb{R}^{W\times H\times C_{y}}$, we firstly extract image features $f(\boldsymbol{x})$ and $f(\boldsymbol{y})$ by shared parallel backbone to save memory and computing resources. Then these features are added in channel dimension to obtain the final $L$ multimodal image features $\boldsymbol{o_{1}},...,\boldsymbol{o_{l}},...,\boldsymbol{o_{L}}\in\mathbb{R}^{W_{1}\times H_{1}\times C_{1}}$. The parameters in $f(\cdot)$ are shared with the OD network.
Although $f(\boldsymbol{x})$ and $f(\boldsymbol{y})$ can describe the characteristics of visible and infrared modalities, they lack insights into the multi-granularity perspective. Therefore, we utilize the branches including the one pixel feature mining module (PFMM) and $L$ region feature refine module (RFRM) to mine the multiple granularities from coarse to fine. The PFMM $B_{0}$ is set the same with feature fusion block in the MetaFusion [17] with the input pair $\boldsymbol{x},\boldsymbol{y}$. We set $1,2,…,l,…L$ to denote each region branch. For branch $l$, a CNN $\varphi_{l}(\cdot)$ is firstly utilized to extract the granularity-wise feature $\varphi_{l}(\boldsymbol{o_{l}})\in\mathbb{R}^{W_{2}\times H_{2}\times C_{2}}$ at the region level. A set of learnable region prompts $\boldsymbol{R}_{l}=\left\{\boldsymbol{r}_{l,m}\in\mathbb{R}^{C_{2}}\right\}_{m=1}^{M_{l}}$ are introduced to define the $M_{l}$ different regions of granularity-wise feature, where $r_{l,m}$ denotes the $m^{\mathrm{th}}$ region prompt at branch $l$. Then the feature vector is mapped into the region mask $\boldsymbol{A}_{l}=\left\{\boldsymbol{a}_{l,m}\in\mathbb{R}^{W_{2}\times H_{2}}\right\}_{m=1}^{M_{l}}$ by conducting the dot product between the feature vector and region prompt followed by batch normalization and a ReLU activation:
@@
where $b_{l,m}(\boldsymbol{o_{l}})$ denotes the $m^{\mathrm{th}}$ region representation and $(i,j)$ denotes the spatial location. These region-level representations are further concatenated to form the observation $B_{l}(\boldsymbol{o_{l}})=\left[b_{l,1}(\boldsymbol{o_{l}}),b_{l,2}(\boldsymbol{o_{l}}),\ldots,b_{l,M_{l}}(\boldsymbol{o_{l}})\right]$ of branch $l$. These multi-grained attentions concentrate operation on the spatial location information and the extent of the regions which are similar to the task requirements. The $B_{1}$, $B_{2}$, …, $B_{L}$ are up-sampled to keep the consistent spatial size with the pixel-level fusion features $B_{0}$. Then, the region-level fusion features $B_{1},B_{2},...,B_{L}$ are assembled by addition operation followed by a Convolution with $1\times 1$ kernel + ReLU to reduce the channel numbers. Highlighted region features are extracted by a Convolution with $1\times 1$ kernel + Sigmoid and then injected into pixel-level features by multiplication and addition. Finally, five Convolutions with $3\times 3$ kernel + ReLU layers are constructed to reconstruct the fusion result $u$.
Coarse-to-Fine Diffusion Process. Diffusion models, inspired by nonequilibrium thermodynamics, are a class of likelihood-based models. DiffusionDet [49] is the first neural network model to utilize the diffusion model for object detection, introducing a novel paradigm that achieves promising results compared to traditional object detection models. Combining diffusion simulation with the diffusion and recovery process of the object box, the Coarse-to-Fine Diffusion Process (CFDP) introduces it as an efficient detection head to assist fusion networks to focus more on object areas. CFDP model defines a Markovian chain of diffusion forward process by gradually adding noise to a set of bounding boxes. The forward noise process is defined as: